Assume country A has a GDP of $2bn and country B $2.5bn. If country A grew on average by 10% and country B on average by 2% per year over the past 15 years, how long will it take for A to overtake B? What factors could change this analysis?
Answer #1

Let, it takes n years for A to overtake B.

Then,

2*(1+10%)^n = 2.5*(1+2%)^n

(1.1/1.02)^n = 2.5/2 = 1.25

1.078^n = 1.25

n = log1.25/log1.078

n = 2.97 years or 3 years

So, A will take 2.97 years or 3 years to pass the economy of B.

Factors such as interest rate in the economy, employment status, inflation and money supply are some of the factor that can change the growth rate and the course of analysis.